### v1

# Mitigating Hallucination of Large Vision-Language Models via Dynamic Logits Calibration

Abstract
Large Vision-Language Models (LVLMs) have demonstrated significant advancements in multimodal understanding, yet they are frequently hampered by hallucination-the generation of text that contradicts visual input. Existing training-free decoding strategies exhibit critical limitations, including the use of static constraints that do not adapt to semantic drift during generation, inefficiency stemming from the need for multiple forward passes, and degradation of detail due to overly rigid intervention rules. To overcome these challenges, this paper introduces Dynamic Logits Calibration (DLC), a novel training-free decoding framework designed to dynamically align text generation with visual evidence at inference time. At the decoding phase, DLC step-wise employs CLIP to assess the semantic alignment between the input image and the generated text sequence. Then, the Relative Visual Advantage (RVA) of candidate tokens is evaluated against a dynamically updated contextual baseline, adaptively adjusting output logits to favor tokens that are visually grounded. Furthermore, an adaptive weighting mechanism, informed by a real-time context alignment score, carefully balances the visual guidance while ensuring the overall quality of the textual output. Extensive experiments conducted across diverse benchmarks and various LVLM architectures (such as LLaVA, InstructBLIP, and MiniGPT-4) demonstrate that DLC significantly reduces hallucinations, outperforming current methods while maintaining high inference efficiency by avoiding multiple forward passes. Overall, we present an effective and efficient decoding-time solution to mitigate hallucinations, thereby enhancing the reliability of LVLMs for more practices. Code will be released on Github.
1 Introduction
The field of multimodal artificial intelligence has witnessed remarkable progress through the development of Large Vision-Language Models (LVLMs) [7, 12, 23, 26, 30, 42, 46, 56]. By integrating visual encoders [35, 51] with large language models (LLMs) [2, 10, 40, 41], contemporary LVLMs have successfully unified visual and textual information within a common semantic framework. These models demonstrate impressive capabilities across diverse visual tasks [3, 4, 21, 31], ranging from straightforward Visual Question Answering (VQA) to sophisticated scene reasoning.
Despite their remarkable advances in multimodal comprehension, LVLMs remain plagued by a critical limitation: hallucination, the generation of text contradicting visual input [4, 5, 17, 22, 57]. This phenomenon introduces deceptive inconsistencies, as the model produces plausible but factually unfaithful descriptions of the observed scene. This issue is especially critical in high-stakes domains like medical diagnosis [6, 24] and autonomous systems [8, 34], where such inconsistencies may incur serious real-world consequence, further hindering the practical deployment of LVLMs.
Figure 1: Comparison of decoding methods across three dimensions: Speed, Correctness, and Detailedness using LLaVA-1.5 [30] (max new tokens: 64), with all experiments conducted on a single NVIDIA A100 GPU. The vertical axis represents correctness, while the horizontal axis shows tokens per second (TPS). Bubble size corresponds to Detailedness of responses. Our method achieves superior accuracy and detailedness while maintaining decoding efficiency.
To address the hallucination issue, numerous training-free decoding strategies have emerged. These approaches have garnered significant attention due to their independence from high-quality annotated datasets and resource-intensive computational requirements. Based on their intervention mechanisms, existing methods can be systematically categorized into three main paradigms: 1) Contrastive Decoding [18, 22, 43]; 2) Guided Decoding Frameworks [9, 13, 53]; and 3) Latent Space Intervention Methods [17, 32, 39, 57]. While these strategies demonstrate some effectiveness in generation tasks, systematic analysis reveals three fundamental limitations. First, static constraint mechanisms fail to dynamically adapt to semantic drift during generation, leaving the model vulnerable to hallucinations in later stages as linguistic priors dominate over visual grounding. Second, most methods require multiple forward passes or regeneration, drastically reducing inference efficiency. Third, rigid intervention rules lack contextual sensitivity, either enforcing overly strict visual alignment or allowing unchecked deviations, thereby distorting the balance between fidelity and creativity.
Addressing these limitations-namely the lack of dynamic adaptation to context, the high inference cost, and the distortion of reality due to rigid rules-demands a deeper investigation. Our work identifies that a core issue underlying hallucination is the increasing reliance on linguistic priors during decoding, causing progressive detachment from visual evidence despite initial alignment. Motivated by this observation and the need to overcome the shortcomings of existing methods, we introduce Dynamic Logits Calibration (DLC) as an efficient and training-free decoding strategy. DLC employs the CLIP model as a dynamic probe to continuously monitor semantic correspondence between the image and generated text during token generation. It evaluates the semantic alignment of both the current text prefix and potential next tokens relative to the image, subsequently adjusting the logits to favor visually grounded options. This mechanism empowers tokens possessing a high "relative visual advantage" to supersede strong linguistic priors, thus diminishing the likelihood of visual hallucination while preserving generation coherence and maintaining high inference speed. As initially evidenced in Figure 1, DLC exhibits superior efficiency compared to other leading approaches. The concerned efficacy is validated through extensive experiments across multiple benchmarks, demonstrating its successful integration into diverse LVLM architectures such as LLaVA [30], InstructBLIP [12], and MiniGPT-4 [56]. In summary, our key contributions are threefold:
  • We revisit the problem of semantic priors in LVLM decoding, highlighting the resulting detachment from visual evidence and onset of hallucination during later generation stages, and revealing a sharp drop in visual alignment for the high-probability tokens selected by the model during hallucination.
  • To address the problem, we craft DLC, a novel training-free scheme, which dynamically adjusts token logits by leveraging CLIP to assess vision-language alignment in real-time, thereby mitigating hallucinations based on contextual visual feedback. The integration of semantic alignment into the logits space offers a new avenue for tackling dynamic vision-language alignment challenges.
  • Comprehensive evaluations on diverse benchmarks confirm the efficacy of DLC, showing significant improvements in reducing hallucinations with appealing running efficiency.
2 Related Work
Hallucination issue in LVLMs. The root cause of LVLM hallucination lies in the inherent tension between parametric knowledge dominance and visual evidence grounding: while users expect models to generate responses strictly conditioned on visual inputs, LVLMs frequently default to priors learned from their textual training corpus, favoring statistically likely but visually ungrounded outputs [15, 16, 22]. This over-reliance on linguistic patterns on top of pixel-accurate reasoning leads to hallucinations when the model’s generative biases override fidelity to the input image (see Appendix A.1 for additional background on LVLMs).
The prior approaches broadly fall into three paradigms. 1) Data-driven optimization [5, 19, 29, 45, 48, 49, 50, 54], including fine-grained contrastive learning, reinforcement learning, and curated instruction tuning with noisy or adversarial datasets. While striving for alignment enhancement, these methods demand expensively annotated data and often risk overfitting issue. 2) Post-hoc correction mechanisms [37, 44, 47, 52, 55], such as auxiliary reviser networks, image regeneration feedback, or external knowledge grounding. Though effective, they introduce significant inference latency and reliance on auxiliaries. 3) Decoding strategy refinement [9, 13, 17, 18, 22, 32, 39, 43, 53, 57], exemplified by VCD [22] that attenuates textual bias by contrasting original and distorted visual inputs, OPERA [17] that utilizes over-trust penalty and a retrospection allocation to alleviate aggregation patterns, and DoLa [11] that mitigates factual hallucinations via layer-wise logit calibration. Despite data dependency being avoided, these training-free methods exhibit critical limitations as discussed in the Introduction. DLC addresses these issues by employing real-time visual alignment assessment (Section 4.1) and adaptive logit modulation (Section 4.2), providing a more faithful decoding stage.
3 Preliminary and Motivation
3.1 Generative Process of LVLMs
An LVLM, parameterized by $\theta$, processes jointly visual images $v$ and textual queries $x$. Typically, a pre-trained network encodes $v$ into visual features. An alignment module (e.g., Q-Former [26] or linear projection [30]) then maps these features into the textual semantic space. These aligned visual representations are concatenated with the embedded text query $x$ to form a joint input representation. Conditioned on this unified input, the LVLM auto-regressively predicts the next token $y_{t}$:
$y_{t}\sim p_{\theta}\left(y_{t}\mid v,x,y_{<t}\right)\propto\exp\left({logit}_{\theta}\left(y_{t}\mid v,x,y_{<t}\right)\right)$ (1)
where $y_{<t}$ are previous tokens and $logit_{\theta}$ is the logits function. Decoding strategies then select the next token, which is appended to the input for subsequent iterations until generation is complete.
Our proposed DLC enhances this generative process as a training-free decoding strategy. Notably, it is model-agnostic, requiring no architectural changes or fine-tuning, allowing flexible application across diverse LVLMs. Furthermore, DLC is compatible with mainstream decoding strategies like nucleus and temperature sampling, seamlessly integrating by modulating logits before token selection. Critically, this approach maintains high inference efficiency, introducing only minimal computational overhead and latency, making it a practical enhancement for visually grounded generation.
3.2 CLIP Model and CLIP Score
Figure 2: Distribution of hallucinatory words by relative position in generated captions, analyzed using the CHAIR benchmark, comparing various decoding methods (ICD, VCD, SID). All methods use sampling decoding strategy, with a Max Token setting of 512. Hallucinations predominantly emerge in the later stages (60%-80%) of generation, indicating progressive semantic drift.
The Contrastive Language-Image Pre-training (CLIP) model [35], trained on extensive datasets of image-text pairs using a contrastive learning objective, learns to embed visual and textual information into a shared, high-dimensional semantic space. Formally, for a given image $v$ and text sequence $s$, we calculate their semantic similarity using cosine similarity, denoted as $\mathrm{CLIP}(v,s)$:
$\mathrm{CLIP}(v,s)=\mathrm{Cosine\_Similarity}(E_{I}(v),E_{T}(s))$ (2)
where $E_{I}(v)$ and $E_{T}(s)$ are embeddings from CLIP’s pre-trained image ($E_{I}$) and text ($E_{T}$) encoders.
A higher $\mathrm{CLIP}(v,s)$ value signifies greater semantic relevance. Section 4 details how our method leverages this score to guide text generation by assessing visual consistency during the decoding.
3.3 Motivation
To understand the origins of hallucinations in LVLMs and identify real-time signals for their mitigation, we empirically analyze the dynamic evolution of the autoregressive generation. As illustrated in Figure 2, a common empirical observation across state-of-the-art decoding methods is that while models often successfully anchor their initial generation in visual evidence, hallucinations tend to emerge progressively, exhibiting a pronounced concentration in the later stages of sequence generation (typically peaking within the 60%-80% range). This pattern of semantic drift-wherein the model increasingly relies on learned linguistic priors and patterns over continuous visual grounding-highlights a tendency to gradually deviate from the visual input as the sequence unfolds. Consequently, this necessitates a closer examination of how visual consistency evolves at the token level during decoding, which is crucial for identifying real-time signals to monitor fidelity and counteract drift.
Building upon this, to dissect the dynamics of semantic drift at a granular level, we leverage the CLIP score (Eq. 2) as a proxy for the semantic alignment between the input image $v$ and the evolving text $y_{<t}$ during decoding. Our subsequent analysis, exemplified by the case study in Figure 3, focuses on the trajectory of visual consistency scores. This approach allows us to contrast the dynamics observed during phases of visually faithful description versus those containing hallucinations.
Figure 3: Visualization of semantic drift and token selection in LVLMs. The central plot tracks Current Contextual Token Alignment (CCTA) against a Historical Baseline $\bar{B}_{t}$, contrasting visually faithful (green) and hallucination (red) phases. Surrounding bar charts show Language Model Logits, Isolated Token Alignment (ITA), and potential Current Contextual Token Alignment (CCTA) if selected for candidate tokens at chosen steps, illustrating the divergence between high logits and low visual alignment during hallucination.
  • Analyzing the Trajectory of Visual Consistency: The central plot in Figure 3 illustrates the evolving visual alignment during text generation. The Current Contextual Token Alignment (CCTA) trajectory quantifies the visual alignment at each step $t$. This score is derived from the CLIP similarity between the input image $v$ and the sequence segment formed by the most recently chosen token $y_{t}$ appended to its immediate preceding context $y_{<t}$ (detailed in Section 4). It effectively measures how well the current state of generation, including the latest token, aligns with the visual input. Complementing this, the Historical Baseline $\bar{B}_{t}$ provides a smoothed average of these CCTA scores from the recent generation history (detailed in Section 4). This baseline acts as a dynamic reference, reflecting the level of visual coherence that the model has recently established and against which the alignment of new tokens can be compared. During the visually-grounded stages of generation, the CCTA trajectory typically remains high, often rising with or closely tracking the $\bar{B}_{t}$, indicating successful selection of tokens that maintain or enhance semantic coherence. However, as generation progresses into hallucination phases, a characteristic downward trend emerges: the CCTA often drops significantly and exhibits increased volatility, frequently falling below the $\bar{B}_{t}$ and other indicative thresholds (e.g., Baseline Threshold). This decline signals a weakening connection to visual evidence and the onset of semantic drift.
  • Examining Candidate Token Selection: The bar charts surrounding the central graph in Figure 3 provide snapshots at different generation stages, illustrating the relationship between the language model’s preference for a token (Logits Score) and that token’s visual consistency. In visually faithful stages, alignment is typical: tokens with high logits scores also tend to exhibit strong visual consistency. Conversely, during hallucinatory stages, a critical divergence is often observed. The model may assign high logits scores to, and consequently select, tokens that possess comparatively low visual consistency scores. This indicates that the selected token is visually inappropriate, either when assessed for its intrinsic alignment with the image as an isolated entity (Isolated Token Alignment (ITA), i.e., $\mathrm{CLIP}(v,c_{i})$) or in terms of how it integrates with the recent sequence context (i.e., its potential Current Contextual Token Alignment, $\mathrm{CLIP}(v,y_{<t}+c_{i})$). This detrimental selection often occurs even when other candidate tokens $c_{i}$ offering better ITA or CCTA scores are available but were assigned lower initial probabilities by the language model.
Synthesizing observations from Figure 3, our key insight is that hallucinations frequently arise when LVLMs, driven by strong linguistic priors, select tokens with high linguistic coherence but poor visual alignment (either intrinsic or contextual), even when better visual alternatives exist. Standard decoding methods, focused on language model likelihood, are ill-equipped to prevent this. They lack the crucial real-time, fine-grained visual consistency checks needed to systematically assess a candidate token’s true visual appropriateness against the unfolding narrative. This critical gap-the failure to dynamically detect and counteract incipient semantic drift by re-evaluating choices against visual evidence-motivates our Dynamic Logits Calibration (DLC), which is designed to explicitly integrate these real-time visual signals, steering generation towards higher visual fidelity and tackling the shortcomings of existing approaches, as detailed in the Section 4 and Figure 4.
4 Methodology
Building upon the insights from our analysis in Section 3.3, we propose a Dynamic Logits Calibration (DLC) mechanism integrated directly into the decoding process. As illustrated in Figure 4, DLC comprises two key components: Real-time Visual Alignment Assessment (Section 4.1) and Adaptive Logit Modulation (Section 4.2).
Figure 4: Overview of Dynamic Logits Calibration (DLC). Given an input image and prompt, DLC first performs Real-time Visual Alignment Assessment on top-k candidate tokens by calculating Current Contextual Token Alignment (CCTA) and Isolated Token Alignment (ITA) scores relative to a historical baseline ($\bar{B}_{t}$). These scores inform the Adaptive Logit Modulation step, which computes Relative Visual Advantage (RVA) and a dynamic modulation factor ($\lambda_{t}$) to adjust the original logits ($L_{t,i}$), favoring visually grounded tokens.
4.1 Real-time Visual Alignment Assessment
At each decoding step $t$, prior to selecting token $y_{t}$, DLC assesses the visual alignment of candidate subsequent tokens with the input image $v$. This assessment involves establishing a dynamic baseline for visual consistency based on recent generation history and then evaluating each candidate token’s potential visual contribution.
First, to quantify the visual grounding of the preceding context, we introduce the History Contextual Token Alignment (HCTA). This score measures the CLIP similarity between the input image $v$ and a fixed size window of $N$ tokens $y_{t-N:t-1}$ immediately preceding the current step, denoted as:
$S_{t}^{\text{HCTA}}=\mathrm{CLIP}(v,y_{t-N:t-1})$ (3)
$S_{t}^{\text{HCTA}}$ reflects the degree to which this recent textual window aligns with the visual input.
To establish a more stable reference reflecting recent trends in visual consistency and mitigate transient fluctuations, we maintain a Historical Baseline $\bar{B}_{t}$. This baseline is calculated as the smoothed average of the $N$ most recent HCTA scores ($S_{k}^{\text{HCTA}}$ for $k<t$) following a brief warm-up period. $\bar{B}_{t}$ thus represents the established level of visual fidelity against which the visual contribution of subsequent candidate tokens will be compared.
Next, for each of the top-k candidate tokens $\{c_{1},...,c_{k}\}$ proposed by the LVLM for the current decoding step $t$, we evaluate their potential impact on visual alignment using two distinct scores:
  • Current Contextual Token Alignment (CCTA). It evaluates the visual alignment resulting from appending a candidate token $c_{i}$ to the preceding contextual window $y_{t-N:t-1}$, as follows:
    $S_{t,i}^{\text{CCTA}}=\mathrm{CLIP}(v,y_{t-N:t-1}\oplus c_{i})$ (4)
    where $\oplus$ denotes an operation of concatenation.
  • Isolated Token Alignment (ITA). This score measures the intrinsic visual alignment of a candidate token $c_{i}$ with the input image $v$, independent of the preceding textual context. It is calculated as:
    $S_{t,i}^{\text{ITA}}=\mathrm{CLIP}(v,c_{i})$ (5)
    Note that $S_{t,i}^{\text{ITA}}$ assesses the candidate’s standalone visual relevance.
Evaluating candidate tokens based solely on their ITA or CCTA scores can be insufficient. For instance, the preceding context may have drifted from strong visual alignment, potentially skewing the contextual score. Conversely, a token’s true visual contribution might only become evident when considered within the sequence, rather than in isolation.
Therefore, to obtain a more robust and nuanced measure of a candidate token’s visual suitability, we compute a combined score for each candidate by integrating ITA and CCTA:
$S_{t,i}^{\text{COMB}}=\frac{S_{t,i}^{\text{CCTA}}+S_{t,i}^{\text{ITA}}}{2}$ (6)
This $S_{t,i}^{\text{COMB}}$ provides a more comprehensive assessment of each candidate’s visual appropriateness, considering both its individual properties and its potential integration into the sequence.
4.2 Adaptive Logit Modulation
Having established the historical visual alignment $\bar{B}_{t}$ and computed the combined scores $S_{t,i}^{\text{COMB}}$ for each candidate token $c_{i}$, the core technique lies in using these signals to adaptively modulate the original logits $L_{t,i}$ produced by LVLM. This modulation aims to increase the likelihood of selecting visually grounded tokens, especially when semantic drift is detected relative to the historical baseline.
The modulation process involves three key steps:
  1. 1.
    Quantifying Relative Visual Contribution: First, to determine how much a candidate $c_{i}$ improves or degrades visual alignment relative to the established context, we calculate its Relative Visual Alignment (RVA) as follows:
    $\mathrm{RVA}_{t,i}=\frac{S_{t,i}^{\text{COMB}}-\bar{B}_{t}}{1-\bar{B}_{t}}$ (7)
    A positive RVA means that the candidate is more visually aligned than the recent baseline $\bar{B}_{t}$, and vice versa. This metric provides a normalized signal of the candidate’s relative visual merit.
  2. 2.
    Adaptive Intervention Strength: The strength of visual guidance applied shall depend on current state of visual grounding. If the generation is already highly aligned with the image, strong intervention is unnecessary and potentially harmful to fluency. Conversely, if alignment is poor, stronger guidance is needed. We determine this dynamic using a Dynamic Guidance Strength $\lambda_{t}$:
    $\lambda_{t}=\alpha\cdot(1-\bar{B}_{t})^{2}$ (8)
    where $\alpha$ controls the maximum modulation strength. The factor $(1-\bar{B}_{t})$ ensures that intervention is applied significantly only when the baseline alignment $\bar{B}_{t}$ drops, thereby minimizing modulation when visual grounding is already sufficient.
  3. 3.
    Applying the Logit Adjustment: Finally, we combine RVA and $\lambda_{t}$ to modify the original logit $L_{t,i}$ for each valid candidate $c_{i}$. The Adjusted Logit $L_{t,i}^{{}^{\prime}}$ is computed multiplicatively:
    $L^{{}^{\prime}}_{t,i}=L_{t,i}\cdot\exp(\lambda_{t}\cdot\sigma(\mathrm{RVA}_{t,i}))$ (9)
    where $\sigma(\cdot)$ denotes the sigmoid function.
This adaptive logit modulation mechanism directly integrates visual feedback into token selection. By dynamically amplifying the probability of visually consistent candidates based on both their relative merit (RVA) and the overall context alignment $\lambda_{t}$, it effectively counteracts misleading semantic priors and promotes the generation of text faithful to the visual input. The complete algorithm and detailed implementation considerations are provided in Appendix A.2 and Appendix A.3, respectively.
5 Experiments
5.1 Experimental Settings
Evaluated LVLMs. We evaluate the effectiveness of our proposed DLC on three state-of-the-art LVLMs: LLaVA-1.5 [30], InstructBLIP [12] and MiniGPT-4 [56]. All benchmarked models adopt a dual-modality architecture, incorporating pre-trained vision encoders like CLIP [35] vision encoder and language models (e.g., Vicuna [10], LLaMA [41]) as fundamental components. Notably, all of models used in our paper are equipped with a 7B LLMs.
Benchmarks.To comprehensively evaluate our proposed method, we employed a suite of established benchmarks. These include:(1) CHAIR [36], for quantifying object hallucination in image captions;(2) GPT-4 Assisted Evaluations [54], to detect more fine-grained hallucination types;(3) GPT-4o assisted evaluations [47], for a comprehensive analysis of hallucinations and text quality; (4) LLaVA-Bench-in-the-Wild [30], on which we performed targeted case studies to demonstrate the efficacy of our proposed DLC in addressing complex tasks and adapting to novel domains. Detailed descriptions of each benchmark, are provided in Appendix A.4.
Baselines. To evaluate our training-free LVLM decoding strategies, we have compared them with standard decoding methods, including Nucleus Sampling, Greedy Sampling. We also assessed the performance of our method against several state-of-the-art decoding approaches: VCD [22], ICD [43], SID [18], and OPERA [17]. For comprehensive comparisons, we implemented VCD, ICD, SID, and DLC using both nucleus sampling (Top-p=1). The reported performance of these baselines is based on our re-implementation using their publicly available code.
5.2 Experimental Results
Table 1: CHAIR hallucination evaluation results. This table evaluates model hallucination using the CHAIRS and CHAIRI metrics, where lower values indicate reduced hallucination and improved performance. TPS indicates tokens per second during generation. All methods use nucleus sampling with Top-p=1. Bold and underlined values indicate the best and second-best results, respectively.
Max Token Methods LLaVA-1.5 InstuctBLIP MiniGPT-4
CS $\downarrow$ CI $\downarrow$ TPS $\uparrow$ CS $\downarrow$ CI $\downarrow$ TPS $\uparrow$ CS $\downarrow$ CI $\downarrow$ TPS $\uparrow$
64 Sample 25.2 9.11 17.4 31.3 11.3 21.9 25.4 9.74 20.0
VCD 24.9 8.15 8.92 27.8 9.56 10.7 22.5 8.27 10.2
ICD 24.0 8.12 8.83 29.3 9.63 9.92 22.6 8.17 9.78
SID 21.9 6.84 8.82 27.7 9.14 9.86 22.5 8.17 9.14
DLC(Ours) 20.9 6.22 16.1 25.4 8.38 13.08 21.6 7.62 8.32
512 Sample 52.3 16.4 17.4 56.2 17.6 21.1 34.8 12.3 19.4
VCD 56.5 17.0 8.74 62.0 18.1 10.4 30.7 10.2 10.0
ICD 52.1 15.4 8.21 65.4 19.7 10.1 32.5 10.5 9.56
SID 51.1 14.4 8.61 58.9 16.8 9.88 30.7 10.2 8.92
DLC(Ours) 38.4 10.8 14.7 51.8 15.2 12.7 29.4 9.62 10.8
CHAIR Evaluation. Following [17, 18], we randomly sample 500 images from COCO [28] validation and prompt the evaluated models with "Please describe this image in detail" to generate descriptions under two maximum new token limits. Moreover, to address CHAIR’s [36] image sensitivity, three distinct image sets were used for robust evaluation. Table 1 shows our method generally surpasses others, validating its efficacy. Notably, our method also exhibits superior generation speed over other elaborately designed approaches. Moreover, as our approach mitigates semantic drift, its superiority is particularly evident with a 512 Max Token setting.
GPT-4 Assisted Evaluation. Following [17, 18, 54], we evaluate on 200 images from the VG dataset using the prompt "Please describe this image in detail," with a maximum of 512 new tokens generated. Additionally, we evaluated the model’s generation speed using tokens per second (TPS). As illustrated in Figure 5, we observe that our DLC achieves state-of-the-art results on metrics related to both hallucinated sentences and words. This indicates that DLC can effectively assist models in mitigating hallucination issues stemming from strong semantic priors. Consistent with observations for OPERA [17], we also find that for certain models, DLC may reduce the length of the generated text, which can be attributed to the reduction of superfluous hallucinatory content.
GPT-4o Assisted Evaluation. To further analyze the hallucinations and text quality for open-end generation tasks, we employ GPT-4o as an advanced evaluation proxy. For evaluation, models generate detailed descriptions for 500 randomly selected COCO [28] images, prompted with: "Please describe this image in detail.". Comparative analyses are performed, contrasting our DLC with several established baselines, including Nucleus Sampling, VCD [22], OPERA [17], and SID [18]. The empirical results, summarized in Table 3, indicate that DLC consistently outperforms these baselines on most metrics. In particular, DLC demonstrates a significant improvement in correctness over conventional sampling-based decoding, while maintaining a comparable degree of detail. The close alignment of GPT-4o’s perceptual and reasoning faculties with human judgment lends substantial credibility to these findings, suggesting that the observed enhancements in hallucination reduction and text quality are indicative of genuine improvements from a human evaluative perspective (full GPT-4/GPT-4o prompt is provided in the Appendix A.9).
Figure 5: GPT-4 assisted benchmark [54]. Six aspects of values are analyzed, including the number of sentences per image (SPI), the tokens per second (TPS), the number of hallucinated sentences per image (HSPI), the number of hallucinated words per image (HWPI), the ratio of hallucinated sentences (HSR), and the ratio of hallucinated words (HWR). Note that larger SPI and TPS, smaller HSPI, HWPI, HSR and HWR are better. Larger radar indicates better performance.
Figure 6: Qualitative comparison on LLaVA-Bench-in-the-Wild [30]. Under identical sampling decoding strategy, our DLC approach is shown to significantly reduce hallucinatory content and enhance descriptive detail compared to baseline methods like Nucleus Sampling and VCD.
Table 2: Assessing hallucination on COCO [28] with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
Settings LLaVA-1.5 InstructBLIP MiniGPT-4
$C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$
Nucleus 5.34 6.18 4.76 5.61 5.15 5.73
Our 7.73 6.87 7.38 6.75 6.64 6.56
VCD 5.55 6.38 5.32 6.09 5.55 5.97
Our 7.45 6.67 6.94 6.54 6.39 6.44
OPERA 6.16 6.59 6.02 6.17 5.98 5.93
Our 7.10 6.67 6.51 6.55 6.16 6.45
SID 5.69 6.54 5.34 6.07 5.55 6.05
Our 7.17 6.55 6.89 6.62 6.54 6.45
Table 3: Assessing hallucination on LLaVA-Bench [30] with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
Settings LLaVA-1.5 InstructBLIP MiniGPT-4
$C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$
Nucleus 4.71 5.79 4.21 5.00 5.17 5.33
Our 8.42 7.04 7.25 6.42 6.00 5.75
VCD 5.25 5.96 4.88 5.54 5.00 5.08
Our 7.46 6.58 6.58 6.17 5.67 5.46
OPERA 5.75 6.75 5.88 5.62 5.75 6.08
Our 7.54 6.88 5.96 6.21 5.79 5.62
SID 5.58 6.79 4.42 5.25 5.38 5.42
Our 6.71 6.08 6.38 6.29 5.88 5.79
Case Study on LLaVA-Bench. To complement our use of dedicated metrics like CHAIR [36] and evaluations facilitated by GPT-4/GPT-4o, we present a qualitative analysis of DLC’s hallucination mitigation capabilities through case studies from the LLaVA-Bench-in-the-Wild dataset [30]. As illustrated in Figure 6 (with additional cases in Appendix A.8), the DLC approach markedly reduces hallucinatory content and enhances descriptive detail under identical image and prompt conditions. These qualitative observations are further supported by a comprehensive GPT-4o evaluation of all LLaVA-Bench generation outputs, detailed in Table 3, which substantiates DLC’s superior performance over competing methods.
Figure 7: Effectiveness of the adaptive factor $(1-\bar{B_{t}})^{2}$ in $\lambda_{t}$.
5.3 Ablation Studies
We primarily validate the effectiveness of the adaptive factor in Dynamic Guidance Strength $\lambda_{t}$. Furthermore, in the Appendix, we explore the influence of the Modulation Strength $\alpha$, the impact of different Window Sizes $N$ (Appendix A.5), and the integration of DLC with various decoding methods (Appendix A.6).
Effectiveness of the adaptive factor in $\lambda_{t}$. To demonstrate the effectiveness of the adaptive factor $(1-\bar{B}_{t})^{2}$ in dynamic guidance strength $\lambda_{t}$, as formulated in Eq. 8, we configure models with and without this factor to generate responses on the COCO [28] dataset. The correctness ($C$) and detailedness ($D$) of these responses are evaluated using GPT-4o. As illustrated in Figure 7, models including this factor achieve higher scores in both accuracy and details compared to their counterparts without the factor, thereby validating its efficacy.
6 Conclusion
This paper introduces Dynamic Logits Calibration (DLC), an innovative training-free decoding framework designed to mitigate hallucinations in LVLMs by dynamically aligning text generation with visual evidence during inference. Observing that hallucinations often arise from the models’ increasing reliance on linguistic priors and progressive detachment from visual input, DLC employs CLIP to step-wise assess image-text semantic alignment and evaluates the Relative Visual Advantage (RVA) of candidate tokens against a dynamically updated contextual baseline. This adaptively adjusts output logits to favor visually grounded tokens, further refined by an adaptive weighting mechanism that balances visual guidance with textual quality. Extensive experiments on various LVLMs demonstrate that DLC significantly reduces hallucinations and outperforms existing methods while maintaining high inference efficiency. In Appendix A.7, we discuss the limitations of our method and its broader implications.
Appendix A Appendix
This appendix provides supplementary materials to support the main findings of our paper. We begin with an overview of Large Vision-Language Models (LVLMs) in Section A.1, offering additional context to help readers understand the foundation of our research. Section A.2 presents a detailed pseudo-code of our proposed method, providing an in-depth algorithmic understanding. Section A.3 elaborates on the implementation details, including parameter settings and computational workflows. For a comprehensive evaluation of our approach, Section A.4 provides further details on the benchmarks used in our experiments. Section A.5 analyzes the impact of two key hyperparameters on performance: the modulation strength $\alpha$ and the historical context window size $N$. Section A.6 demonstrates the integration of DLC with various decoding strategies, including beam search, greedy search, top-k sampling, and temperature scaling. Section A.7 discusses the inherent limitations of our method, while Section A.8 presents additional comparative results of our approach against other methods on the LLaVA-Bench benchmark. Finally, Section A.9 showcases the prompts used for GPT-4 and GPT-4o evaluations.
A.1 More backgrounds about LVLMs
The development of LVLMs has evolved from BERT-based [14, 33, 38] cross-modal alignment to LLM-powered [1, 7, 12, 26, 30, 56] architectures. Building upon foundational text-image alignment techniques like CLIP [35] and BLIP [25], contemporary LVLMs integrate visual encoders with LLM decoders through distinct optimization strategies. For instance, LLaVA [30] employs lightweight adapters combined with instruction tuning, while InstructBLIP [12] and MiniGPT-4 [56] utilize Q-Former modules for visual token compression. Despite these innovations, significant challenges persist, particularly severe hallucination issues arising from misaligned multimodal representations. To systematically analyze mitigation strategies, we have examined with three representative architectures, including LLaVA, InstructBLIP, and MiniGPT-4, whose results demonstrate that various technical approaches commonly share critical vulnerabilities to the hallucination issue.
A.2 Pseudo-code of Our Method
In this section, we provide a detailed pseudo-code of our Dynamic Logits Calibration (DLC) method, as shown in Algorithm 1. The pseudo-code encompasses both key modules: Real-time Visual Alignment Assessment and Adaptive Logit Modulation.
Algorithm 1 Dynamic Logits Calibration (DLC)
1:Input image $v$, text prompt $x$, vision-language model $M_{\theta}$, CLIP model, window size $N$, modulation strength $\alpha$
2:Generated text sequence $y$
3:$y\leftarrow x$triangleright Initialize output sequence with input prompt
4:$\bar{B}\leftarrow 0$triangleright Initialize historical baseline
[…]
7:     $L_{t}\leftarrow M_{\theta}(y_{<t},v)$triangleright Get original logits for current step
8:     if $t>|x|+3$ then $\triangleright$ Start applying DLC after warm-up period
9:         $y_{t-N:t-1}\leftarrow$ get previous $N$ tokens $\triangleright$ Get preceding context window
10:         $S_{t}^{\text{HCTA}}\leftarrow\text{CLIP}(v,y_{t-N:t-1})$triangleright Calculate History Contextual Token Alignment
11:         Add $S_{t}^{\text{HCTA}}$ to HCTA_history
12:         if HCTA_history length $>N$ then
13:              Remove oldest element from HCTA_history
14:         end if
[…]
16:         $C\leftarrow\text{get top-k candidate tokens}(L_{t})$triangleright Get top-k candidates
17:         for each candidate token $c_{i}\in C$ do
18:              $S_{t,i}^{\text{CCTA}}\leftarrow\text{CLIP}(v,y_{t-N:t-1}\oplus c_{i})$triangleright Current Contextual Token Alignment
19:              $S_{t,i}^{\text{ITA}}\leftarrow\text{CLIP}(v,c_{i})$triangleright Isolated Token Alignment
20:              $S_{t,i}^{\text{COMB}}\leftarrow\frac{1}{2}(S_{t,i}^{\text{CCTA}}+S_{t,i}^{\text{ITA}})$triangleright Combined score
21:              $\text{RVA}_{t,i}\leftarrow\frac{S_{t,i}^{\text{COMB}}-\bar{B}_{t}}{1-\bar{B}_{t}}$triangleright Relative Visual Advantage
22:              $\lambda_{t}\leftarrow\alpha\cdot(1-\bar{B}_{t})^{2}$triangleright Dynamic guidance strength
23:              $L^{\prime}_{t,i}\leftarrow L_{t,i}\cdot\exp(\lambda_{t}\cdot\sigma(\text{RVA}_{t,i}))$triangleright Apply logit adjustment
24:         end for
25:         Replace original logits $L_{t}$ with adjusted logits $L^{\prime}_{t}$
26:     end if
27:     $y_{t}\leftarrow\text{sample}(L_{t})$triangleright Sample next token based on adjusted logits
28:     $y\leftarrow y\oplus y_{t}$triangleright Append new token to output sequence
29:     if $y_{t}$ is end token then
30:         break
31:     end if
32:end for
33:return $y$
As shown in Algorithm 1, DLC dynamically adjusts the output logits of LVLMs during decoding to mitigate hallucinations. First, the model undergoes a brief warm-up period (line 6), allowing generation of sufficient context for CLIP assessment. Subsequently, at each decoding step, the current History Contextual Token Alignment (HCTA) score is calculated (lines 7-8), and a historical baseline $\bar{B}_{t}$ is maintained (lines 9-13).
For each candidate token, DLC computes three key scores:
  • Current Contextual Token Alignment (CCTA): measuring the visual alignment of the candidate token with the preceding context (line 16)
  • Isolated Token Alignment (ITA): measuring the inherent alignment of the individual candidate token with the input image (line 17)
  • Combined score: integrating CCTA and ITA by averaging (line 18)
Based on these scores, DLC calculates the Relative Visual Advantage (RVA) for each candidate token (line 19), which represents the improvement in visual alignment relative to the historical baseline. The dynamic guidance strength $\lambda_{t}$ (line 20) adaptively adjusts according to the current state of visual alignment, providing stronger intervention when alignment decreases.
Finally, DLC modifies the original logits through RVA and $\lambda_{t}$ (line 21), thereby enhancing the probability of visually consistent tokens. The entire process integrates into the normal decoding flow without requiring additional forward passes or fine-tuning, thus providing a computationally efficient and effective solution for hallucination mitigation.
A.3 Implementation Details.
In this section, we provide comprehensive implementation details of our Dynamic Logits Calibration (DLC) method, including model configuration, parameter settings, and computational workflows. These details are essential for reproducing our results and understanding the practical aspects of DLC.
We compute all visual alignment scores using the SigLIP model [51]. The historical baseline $\bar{B}_{t}$ is calculated over a sliding window of $N=8$ past context scores, following a 3-step warm-up period. For the top-k candidates $(k=50)$ at each step, the combined score $S_{t,i}^{\text{COMB}}$ is computed. The adaptive modulation employs a maximum intervention strength $\alpha=3$, and the standard sigmoid function $\sigma(\cdot)$ for logit adjustment. Our experiments were run on a single A100 GPU.
A.4 Further Details on Benchmarks
To comprehensively evaluate our DLC method, we utilized several benchmarks specifically designed to assess hallucination in vision-language models. This section provides additional details about these benchmarks, including dataset characteristics, evaluation protocols, and metrics interpretation.
(1) CHAIR [36] quantifies object hallucination in image captions by measuring the ratio of objects mentioned in captions but absent from ground-truth labels. It includes two variants: CHAIRI ($C_{I}$), calculating hallucinated objects versus all generated objects, and CHAIRS ($C_{S}$), assessing sentences with hallucinations versus total sentences. These two metrics are defined as follows::
$C_{I}=\frac{|\{\text{hallucinated objects}\}|}{|\{\text{all mentioned objects}\}|},C_{S}=\frac{|\{\text{captions with hallucinated objects}\}|}{|\{\text{allcaptions}\}|}$ (10)
Lower scores indicate better hallucination suppression.
(2) GPT-4 Assisted Evaluations [54] function as a crucial complement to conventional assessment methodologies. While traditional metrics such as CHAIR effectively quantify hallucinations at the object-existence level, they demonstrate inadequacy in detecting more fine-grained hallucination types, including positional, relational, and attribute-based inaccuracies. To address this limitation, GPT-4 assisted evaluation employs detailed object-level descriptions from the Visual Genome (VG) [20] dataset as ground-truth reference, leveraging GPT-4’s sophisticated capabilities for precise hallucination detection and classification. The complete prompt template is provided in the Appendix A.9.
(3) GPT-4o assisted evaluations [^1] We utilize GPT-4o for assisted evaluation, as it offers enhanced capabilities over previous vision-capable models like GPT-4V, which is also being phased out by OpenAI. [47] further conduct a comprehensive analysis of hallucinations and text quality in open-ended generation tasks. Harnessing GPT-4o’s sophisticated capabilities, [47] established a numerical assessment protocol utilizing a 0-10 scale, which evaluates two essential dimensions: Accuracy-the extent to which responses align with image content without fabrication, and Detailedness-the degree of informational richness present in the responses.
(4) LLaVA-Bench-in-the-Wild [30] is designed to evaluate LVLMs’ proficiency in addressing complex tasks and their adaptability to novel domains. We perform targeted case studies on this dataset to demonstrate the efficacy of our proposed DLC.
A.5 Sensitivity Analysis of Key Hyperparameters: $\alpha$ and $N$
The performance of our DLC method is influenced by two key hyperparameters: the modulation strength $\alpha$ and the context window size $N$. In this section, we conduct a comprehensive sensitivity analysis to understand how these parameters affect the model’s hallucination mitigation capability and generation quality. Experiments were conducted using a range of values for these parameters across three distinct LVLMs, with performance assessed by the CHAIR benchmark [36] (max new tokens: 512 for all ablation studies).
Table 4: Impact of Modulation Strengths $\alpha$ on CHAIR Metrics for Different LVLMs and Decoding Strategies
Method $\alpha$ LLaVA-1.5 InstructBLIP MiniGPT-4
$C_{S}\downarrow$ $C_{I}\downarrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $C_{S}\downarrow$ $C_{I}\downarrow$
Greedy Search 1.0 33.8 9.85 47.9 13.2 32.6 10.3
2.0 34.1 10.0 48.1 13.5 33.7 10.1
3.0 34.8 10.1 49.6 13.5 33.0 9.89
4.0 35.3 9.87 50.5 14.2 33.5 10.0
5.0 36.8 10.5 51.0 14.4 33.9 10.6
Nucleus Sampling 1.0 41.8 12.7 55.4 15.4 34.6 11.1
2.0 40.3 11.4 53.2 14.8 34.0 10.6
3.0 38.4 10.9 51.8 15.2 32.7 10.7
4.0 37.5 10.8 50.8 14.7 33.6 10.7
5.0 37.5 10.9 50.6 14.5 33.6 10.9
Table 5: Impact of Window Size $N$ on CHAIR Metrics for Different LVLMs and Decoding Strategies
Method $N$ LLaVA-1.5 InstructBlip MiniGPT4
$C_{S}\downarrow$ $C_{I}\downarrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $C_{S}\downarrow$ $C_{I}\downarrow$
Greedy Search 4 34.4 9.45 52.2 14.6 35.4 11.1
8 35.1 9.72 49.0 12.8 33.4 10.3
12 34.4 9.51 48.0 13.5 33.2 9.85
16 37.0 9.84 47.0 13.2 31.0 9.69
Nucleus Sampling 4 38.6 11.1 51.4 14.3 31.0 10.2
8 39.4 10.9 51.0 14.3 32.0 11.2
12 38.6 11.5 54.8 14.8 32.4 10.1
16 37.6 11.4 50.4 15.0 34.0 12.6
First, we analyze the effect of varying modulation strengths $\alpha$. The optimal setting for $\alpha$ is contingent upon the specific LVLM and the decoding strategy employed. As detailed in Table 4, under Greedy Search, lower $\alpha$ values, particularly $\alpha=1.0$, tend to be more effective for LLaVA-1.5 and InstructBLIP. For MiniGPT-4 under Greedy Search, competitive results are observed at both $\alpha=1.0$ and $\alpha=3.0$. Conversely, when employing Nucleus Sampling, moderate to higher $\alpha$ values (typically ranging from 3.0 to 5.0) generally yield improved CHAIR scores, with LLaVA-1.5 and MiniGPT-4 showing optimal performance around $\alpha=4.0$ and $\alpha=3.0$ respectively, while InstructBLIP benefits from $\alpha$ values up to 5.0. Therefore, while a global value of $\alpha=3.0$ is adopted for the main experiments in this paper, these findings underscore that tailoring $\alpha$ to specific model architectures and decoding methodologies can yield further performance enhancements.
Next, we examine the impact of the window size $N$. The results, presented in Table 5, indicate varied responses based on the model and decoding strategy. Under Greedy Decoding, both InstructBLIP and MiniGPT-4 generally deliver improved performance with larger window sizes. Conversely, for LLaVA-1.5 using Greedy Decoding, the influence of $N$ is less definitive, suggesting that smaller to moderate $N$ values might be more advantageous. Nucleus Sampling exhibits distinct responses to variations in $N$: MiniGPT-4 achieves its best performance with a small window, with larger $N$ values leading to a decline; LLaVA-1.5 shows more fluctuating results; and InstructBLIP, which generally performs suboptimally with Nucleus Sampling, does not yield significant improvements with changes in $N$. Thus, while a uniform window size of $N=8$ is currently utilized in our main experiments, our findings highlight that tailoring $N$ to specific models and decoding methodologies can substantially boost performance.
In summary, both the modulation strength $\alpha$ and the window size $N$ are influential hyperparameters for DLC. While default values are used for general demonstration in the main body of this paper, the ablation studies presented here clearly indicate that optimal performance is achieved when these parameters are carefully tuned according to the specific LVLM and the selected decoding strategy.
A.6 DLC Performance with Various Decoding Strategies
Our proposed DLC demonstrates compatibility with various mainstream decoding strategies. To comprehensively evaluate its effectiveness, in addition to its performance with common strategies such as nucleus sampling and greedy sampling, we conducted further experiments on the LLaVA-1.5 [30] model using the MSCOCO dataset [28] and the CHAIR benchmark [36]. These experiments explored other decoding methods and parameters, specifically: Top-p sampling (p=0.9), Top-k sampling (k=50), and Top-k sampling (k=50) with varying temperatures (t=1.5 and t=0.8). The results, presented in Figure 8, consistently indicate that regardless of the sampling method employed, DLC effectively reduces hallucinations and enhances overall model performance. This substantiates the robustness and effectiveness of our proposed DLC across diverse decoding strategies.
Figure 8: Results of different decoding strategies.
A.7 Discussion of Methodological Limitations and Social impacts.
In this section, we delineate two limitations inherent to the proposed DLC method:
Scope of Applicability: The DLC method is primarily optimized for open-ended, descriptive text generation scenarios, aiming to mitigate semantic drift that may occur during the generation of lengthy texts by dynamically calibrating the probabilities of candidate tokens. Consequently, its advantages may not be fully realized in contexts requiring minimalistic answers, such as "yes/no" judgments (e.g., as encountered in tasks like the POPE benchmark [27]). The core mechanism of DLC relies on the generated textual context for dynamic adjustments; for extremely short answers, such as single "yes" or "no" responses, the lack of sufficient contextual information for effective logits calibration might constrain its performance. To address this limitation, future work could explore transforming such concise answers (e.g., "yes" or "no") into complete declarative sentences. For instance, for object presence verification, sentences like "There is an <object> in the image" versus "There is not an <object> in the image" could be constructed. Subsequently, CLIP could be employed to compute the semantic similarity between these statements and the input image, thereby assisting the model in selecting a more accurate response. This direction warrants further exploration and validation.
Dependency on CLIP Model Performance: The efficacy of DLC is intrinsically linked to the performance of the employed CLIP model. The CLIP model is tasked with assessing the semantic similarity between images and text, which serves as the foundational input for DLC’s logits calibration process. Therefore, if the CLIP model itself exhibits deficiencies or biases in accurately capturing the semantic associations for specific image-text pairs-stemming, for example, from biases in its training data or an insufficient understanding of certain complex, fine-grained visual concepts-this can directly impact the precision and overall effectiveness of the DLC calibration. Theoretically, leveraging more advanced models with superior vision-text alignment capabilities is anticipated to further enhance the performance ceiling of DLC.
Social impacts. Dynamic Logits Calibration (DLC) holds significant promise for advancing the field of LVLMs. By offering a novel training-free decoding framework that dynamically aligns text generation with visual evidence, DLC effectively mitigates the pervasive issue of hallucination in LVLMs. This approach overcomes critical limitations of existing strategies, such as reliance on static constraints and the inefficiency of multiple forward passes, thereby enhancing the reliability and practical applicability of these models without incurring additional training costs. DLC not only provides an effective and efficient solution to a key challenge but also serves as an inspiration for the research community to explore new pathways for improving visual grounding and tackling dynamic vision-language alignment, fostering the development of more robust and trustworthy LVLMs.
A.8 Extended Comparative Results on LLaVA-Bench
In this section, we present additional comparative results of our DLC approach against other methods on the LLaVA-Bench [30]. These extended evaluations provide a more comprehensive understanding of our method’s performance across various question types and scenarios.
Figure 9: Case Study: LLaVA-1.5 on LLaVA-Bench-in-the-Wild (Hallucinations Marked Red).
A.9 Prompts for GPT-4 and GPT-4o Evaluation
Figure 10: Prompts of GPT-4 for evaluations.
Figure 11: Prompts of GPT-4o for evaluations.

### v2

@@
 1 Introduction
-The field of multimodal artificial intelligence has witnessed remarkable progress through the development of Large Vision-Language Models (LVLMs) [7, 12, 23, 26, 30, 42, 46, 56].
-By integrating visual encoders [35, 51] with large language models (LLMs) [2, 10, 40, 41], contemporary LVLMs have successfully unified visual and textual information within a common semantic framework.
-These models demonstrate impressive capabilities across diverse visual tasks [3, 4, 21, 31], ranging from straightforward Visual Question Answering (VQA) to sophisticated scene reasoning.
-Despite their remarkable advances in multimodal comprehension, LVLMs remain plagued by a critical limitation: hallucination, the generation of text contradicting visual input [4, 5, 17, 22, 57].
+The field of multimodal artificial intelligence has witnessed remarkable progress through the development of Large Vision-Language Models (LVLMs) (Chen et al., 2023; Dai et al., 2023; Li et al., 2024, 2023b; Liu et al., 2023; Wang et al., 2024a; Ye et al., 2024; Zhu et al., 2023).
+By integrating visual encoders (Radford et al., 2021; Zhai et al., 2023) with large language models (LLMs) (Bai et al., 2023a; Chiang et al., 2023; Touvron et al., 2023a, b), contemporary LVLMs have successfully unified visual and textual information within a common semantic framework.
+These models demonstrate impressive capabilities across diverse visual tasks (Bai et al., 2023b, 2024; Lai et al., 2024; Liu et al., 2024b), ranging from straightforward Visual Question Answering (VQA) to sophisticated scene reasoning.
+Despite their remarkable advances in multimodal comprehension, LVLMs remain plagued by a critical limitation: hallucination, the generation of text contradicting visual input (Bai et al., 2024; Chen et al., 2025; Huang et al., 2024; Leng et al., 2024; Zhuang et al., 2025).
 This phenomenon introduces deceptive inconsistencies, as the model produces plausible but factually unfaithful descriptions of the observed scene.
-This issue is especially critical in high-stakes domains like medical diagnosis [6, 24] and autonomous systems [8, 34], where such inconsistencies may incur serious real-world consequence, further hindering the practical deployment of LVLMs.
-Figure 1: Comparison of decoding methods across three dimensions: Speed, Correctness, and Detailedness using LLaVA-1.5 [30] (max new tokens: 64), with all experiments conducted on a single NVIDIA A100 GPU.
+This issue is especially critical in high-stakes domains like medical diagnosis (Chen et al., 2024a; Li et al., 2023a) and autonomous systems (Chen et al., 2024b; Nie et al., 2024), where such inconsistencies may incur serious real-world consequence, further hindering the practical deployment of LVLMs.
+Figure 1: Comparison of decoding methods across three dimensions: Speed, Correctness, and Detailedness using LLaVA-1.5 (Liu et al., 2023) (max new tokens: 64), with all experiments conducted on a single NVIDIA A100 GPU.
 The vertical axis represents correctness, while the horizontal axis shows tokens per second (TPS).
@@
 These approaches have garnered significant attention due to their independence from high-quality annotated datasets and resource-intensive computational requirements.
-Based on their intervention mechanisms, existing methods can be systematically categorized into three main paradigms: 1) Contrastive Decoding [18, 22, 43]; 2) Guided Decoding Frameworks [9, 13, 53]; and 3) Latent Space Intervention Methods [17, 32, 39, 57].
+Based on their intervention mechanisms, existing methods can be systematically categorized into three main paradigms: 1) Contrastive Decoding (Huo et al., 2025; Leng et al., 2024; Wang et al., 2024b); 2) Guided Decoding Frameworks (Chen et al., 2024c; Deng et al., 2024; Zhao et al., 2024); and 3) Latent Space Intervention Methods (Huang et al., 2024; Liu et al., 2025; Tang et al., 2025; Zhuang et al., 2025).
 While these strategies demonstrate some effectiveness in generation tasks, systematic analysis reveals three fundamental limitations.
@@
 As initially evidenced in Figure 1, DLC exhibits superior efficiency compared to other leading approaches.
-The concerned efficacy is validated through extensive experiments across multiple benchmarks, demonstrating its successful integration into diverse LVLM architectures such as LLaVA [30], InstructBLIP [12], and MiniGPT-4 [56].
+The concerned efficacy is validated through extensive experiments across multiple benchmarks, demonstrating its successful integration into diverse LVLM architectures such as LLaVA (Liu et al., 2023), InstructBLIP (Dai et al., 2023), and MiniGPT-4 (Zhu et al., 2023).
 In summary, our key contributions are threefold:
@@
 Hallucination issue in LVLMs.
-The root cause of LVLM hallucination lies in the inherent tension between parametric knowledge dominance and visual evidence grounding: while users expect models to generate responses strictly conditioned on visual inputs, LVLMs frequently default to priors learned from their textual training corpus, favoring statistically likely but visually ungrounded outputs [15, 16, 22].
+The root cause of LVLM hallucination lies in the inherent tension between parametric knowledge dominance and visual evidence grounding: while users expect models to generate responses strictly conditioned on visual inputs, LVLMs frequently default to priors learned from their textual training corpus, favoring statistically likely but visually ungrounded outputs (Gupta et al., 2022; Han et al., 2022; Leng et al., 2024).
 This over-reliance on linguistic patterns on top of pixel-accurate reasoning leads to hallucinations when the model’s generative biases override fidelity to the input image (see Appendix A.1 for additional background on LVLMs).
 The prior approaches broadly fall into three paradigms.
-1) Data-driven optimization [5, 19, 29, 45, 48, 49, 50, 54], including fine-grained contrastive learning, reinforcement learning, and curated instruction tuning with noisy or adversarial datasets.
+1) Data-driven optimization (Chen et al., 2025; Jiang et al., 2024; Liu et al., 2024a; Xiao et al., 2024; Yu et al., 2024a, b; Yue et al., 2024; Zhao et al., 2023), including fine-grained contrastive learning, reinforcement learning, and curated instruction tuning with noisy or adversarial datasets.
 While striving for alignment enhancement, these methods demand expensively annotated data and often risk overfitting issue.
-2) Post-hoc correction mechanisms [37, 44, 47, 52, 55], such as auxiliary reviser networks, image regeneration feedback, or external knowledge grounding.
+2) Post-hoc correction mechanisms (Sahu et al., 2024; Wu et al., 2024; Yin et al., 2024; Zhang et al., 2025; Zhou et al., 2024), such as auxiliary reviser networks, image regeneration feedback, or external knowledge grounding.
 Though effective, they introduce significant inference latency and reliance on auxiliaries.
-3) Decoding strategy refinement [9, 13, 17, 18, 22, 32, 39, 43, 53, 57], exemplified by VCD [22] that attenuates textual bias by contrasting original and distorted visual inputs, OPERA [17] that utilizes over-trust penalty and a retrospection allocation to alleviate aggregation patterns, and DoLa [11] that mitigates factual hallucinations via layer-wise logit calibration.
+3) Decoding strategy refinement (Chen et al., 2024c; Deng et al., 2024; Huang et al., 2024; Huo et al., 2025; Leng et al., 2024; Liu et al., 2025; Tang et al., 2025; Wang et al., 2024b; Zhao et al., 2024; Zhuang et al., 2025), exemplified by VCD (Leng et al., 2024) that attenuates textual bias by contrasting original and distorted visual inputs, OPERA (Huang et al., 2024) that utilizes over-trust penalty and a retrospection allocation to alleviate aggregation patterns, and DoLa (Chuang et al., 2024) that mitigates factual hallucinations via layer-wise logit calibration.
 Despite data dependency being avoided, these training-free methods exhibit critical limitations as discussed in the Introduction.
@@
 Typically, a pre-trained network encodes $v$ into visual features.
-An alignment module (e.g., Q-Former [26] or linear projection [30]) then maps these features into the textual semantic space.
+An alignment module (e.g., Q-Former (Li et al., 2023b) or linear projection (Liu et al., 2023)) then maps these features into the textual semantic space.
 These aligned visual representations are concatenated with the embedded text query $x$ to form a joint input representation.
@@
 Hallucinations predominantly emerge in the later stages (60%-80%) of generation, indicating progressive semantic drift.
-The Contrastive Language-Image Pre-training (CLIP) model [35], trained on extensive datasets of image-text pairs using a contrastive learning objective, learns to embed visual and textual information into a shared, high-dimensional semantic space.
+The Contrastive Language-Image Pre-training (CLIP) model (Radford et al., 2021), trained on extensive datasets of image-text pairs using a contrastive learning objective, learns to embed visual and textual information into a shared, high-dimensional semantic space.
 Formally, for a given image $v$ and text sequence $s$, we calculate their semantic similarity using cosine similarity, denoted as $\mathrm{CLIP}(v,s)$:
@@
 Evaluated LVLMs.
-We evaluate the effectiveness of our proposed DLC on three state-of-the-art LVLMs: LLaVA-1.5 [30], InstructBLIP [12] and MiniGPT-4 [56].
-All benchmarked models adopt a dual-modality architecture, incorporating pre-trained vision encoders like CLIP [35] vision encoder and language models (e.g., Vicuna [10], LLaMA [41]) as fundamental components.
+We evaluate the effectiveness of our proposed DLC on three state-of-the-art LVLMs: LLaVA-1.5 (Liu et al., 2023), InstructBLIP (Dai et al., 2023) and MiniGPT-4 (Zhu et al., 2023).
+All benchmarked models adopt a dual-modality architecture, incorporating pre-trained vision encoders like CLIP (Radford et al., 2021) vision encoder and language models (e.g., Vicuna (Chiang et al., 2023), LLaMA (Touvron et al., 2023b)) as fundamental components.
 Notably, all of models used in our paper are equipped with a 7B LLMs.
 Benchmarks.To comprehensively evaluate our proposed method, we employed a suite of established benchmarks.
-These include:(1) CHAIR [36], for quantifying object hallucination in image captions;(2) GPT-4 Assisted Evaluations [54], to detect more fine-grained hallucination types;(3) GPT-4o assisted evaluations [47], for a comprehensive analysis of hallucinations and text quality; (4) LLaVA-Bench-in-the-Wild [30], on which we performed targeted case studies to demonstrate the efficacy of our proposed DLC in addressing complex tasks and adapting to novel domains.
+These include:(1) CHAIR (Rohrbach et al., 2018), for quantifying object hallucination in image captions;(2) GPT-4 Assisted Evaluations (Zhao et al., 2023), to detect more fine-grained hallucination types;(3) GPT-4o assisted evaluations (Yin et al., 2024), for a comprehensive analysis of hallucinations and text quality; (4) LLaVA-Bench-in-the-Wild (Liu et al., 2023), on which we performed targeted case studies to demonstrate the efficacy of our proposed DLC in addressing complex tasks and adapting to novel domains.
 Detailed descriptions of each benchmark, are provided in Appendix A.4.
@@
 To evaluate our training-free LVLM decoding strategies, we have compared them with standard decoding methods, including Nucleus Sampling, Greedy Sampling.
-We also assessed the performance of our method against several state-of-the-art decoding approaches: VCD [22], ICD [43], SID [18], and OPERA [17].
+We also assessed the performance of our method against several state-of-the-art decoding approaches: VCD (Leng et al., 2024), ICD (Wang et al., 2024b), SID (Huo et al., 2025), and OPERA (Huang et al., 2024).
 For comprehensive comparisons, we implemented VCD, ICD, SID, and DLC using both nucleus sampling (Top-p=1).
@@
 Bold and underlined values indicate the best and second-best results, respectively.
-Max Token Methods LLaVA-1.5 InstuctBLIP MiniGPT-4
+Max
+Token
+Methods LLaVA-1.5 InstuctBLIP MiniGPT-4
 CS $\downarrow$ CI $\downarrow$ TPS $\uparrow$ CS $\downarrow$ CI $\downarrow$ TPS $\uparrow$ CS $\downarrow$ CI $\downarrow$ TPS $\uparrow$
@@
 CHAIR Evaluation.
-Following [17, 18], we randomly sample 500 images from COCO [28] validation and prompt the evaluated models with "Please describe this image in detail" to generate descriptions under two maximum new token limits.
-Moreover, to address CHAIR’s [36] image sensitivity, three distinct image sets were used for robust evaluation.
+Following (Huang et al., 2024; Huo et al., 2025), we randomly sample 500 images from COCO (Lin et al., 2014) validation and prompt the evaluated models with "Please describe this image in detail" to generate descriptions under two maximum new token limits.
+Moreover, to address CHAIR’s (Rohrbach et al., 2018) image sensitivity, three distinct image sets were used for robust evaluation.
 Table 1 shows our method generally surpasses others, validating its efficacy.
@@
 GPT-4 Assisted Evaluation.
-Following [17, 18, 54], we evaluate on 200 images from the VG dataset using the prompt "Please describe this image in detail," with a maximum of 512 new tokens generated.
+Following (Huang et al., 2024; Huo et al., 2025; Zhao et al., 2023), we evaluate on 200 images from the VG dataset using the prompt "Please describe this image in detail," with a maximum of 512 new tokens generated.
 Additionally, we evaluated the model’s generation speed using tokens per second (TPS).
@@
 This indicates that DLC can effectively assist models in mitigating hallucination issues stemming from strong semantic priors.
-Consistent with observations for OPERA [17], we also find that for certain models, DLC may reduce the length of the generated text, which can be attributed to the reduction of superfluous hallucinatory content.
+Consistent with observations for OPERA (Huang et al., 2024), we also find that for certain models, DLC may reduce the length of the generated text, which can be attributed to the reduction of superfluous hallucinatory content.
 GPT-4o Assisted Evaluation.
 To further analyze the hallucinations and text quality for open-end generation tasks, we employ GPT-4o as an advanced evaluation proxy.
-For evaluation, models generate detailed descriptions for 500 randomly selected COCO [28] images, prompted with: "Please describe this image in detail.".
-Comparative analyses are performed, contrasting our DLC with several established baselines, including Nucleus Sampling, VCD [22], OPERA [17], and SID [18].
+For evaluation, models generate detailed descriptions for 500 randomly selected COCO (Lin et al., 2014) images, prompted with: "Please describe this image in detail.".
+Comparative analyses are performed, contrasting our DLC with several established baselines, including Nucleus Sampling, VCD (Leng et al., 2024), OPERA (Huang et al., 2024), and SID (Huo et al., 2025).
 The empirical results, summarized in Table 3, indicate that DLC consistently outperforms these baselines on most metrics.
@@
 The close alignment of GPT-4o’s perceptual and reasoning faculties with human judgment lends substantial credibility to these findings, suggesting that the observed enhancements in hallucination reduction and text quality are indicative of genuine improvements from a human evaluative perspective (full GPT-4/GPT-4o prompt is provided in the Appendix A.12).
-Figure 5: GPT-4 assisted benchmark [54].
+Figure 5: GPT-4 assisted benchmark (Zhao et al., 2023).
 Six aspects of values are analyzed, including the number of sentences per image (SPI), the tokens per second (TPS), the number of hallucinated sentences per image (HSPI), the number of hallucinated words per image (HWPI), the ratio of hallucinated sentences (HSR), and the ratio of hallucinated words (HWR).
@@
 Larger radar indicates better performance.
-Figure 6: Qualitative comparison on LLaVA-Bench-in-the-Wild [30].
+Figure 6: Qualitative comparison on LLaVA-Bench-in-the-Wild (Liu et al., 2023).
 Under identical sampling decoding strategy, our DLC approach is shown to significantly reduce hallucinatory content and enhance descriptive detail compared to baseline methods like Nucleus Sampling and VCD.
-Table 2: Assessing hallucination on COCO [28] with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
+Table 2: Assessing hallucination on COCO (Lin et al., 2014) with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
 Settings LLaVA-1.5 InstructBLIP MiniGPT-4
@@
 Our 7.17 6.55 6.89 6.62 6.54 6.45
-Table 3: Assessing hallucination on LLaVA-Bench [30] with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
+Table 3: Assessing hallucination on LLaVA-Bench (Liu et al., 2023) with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
 Settings LLaVA-1.5 InstructBLIP MiniGPT-4
@@
 Case Study on LLaVA-Bench.
-To complement our use of dedicated metrics like CHAIR [36] and evaluations facilitated by GPT-4/GPT-4o, we present a qualitative analysis of DLC’s hallucination mitigation capabilities through case studies from the LLaVA-Bench-in-the-Wild dataset [30].
+To complement our use of dedicated metrics like CHAIR (Rohrbach et al., 2018) and evaluations facilitated by GPT-4/GPT-4o, we present a qualitative analysis of DLC’s hallucination mitigation capabilities through case studies from the LLaVA-Bench-in-the-Wild dataset (Liu et al., 2023).
 As illustrated in Figure 6 (with additional cases in Appendix A.13), the DLC approach markedly reduces hallucinatory content and enhances descriptive detail under identical image and prompt conditions.
@@
 5.3 Ablation Studies
-We primarily validate the effectiveness of the adaptive factor in Dynamic Guidance Strength $\lambda_{t}$.
-Furthermore, in the Appendix, we explore the influence of the Modulation Strength $\alpha$, the impact of different Window Sizes $N$ (Appendix A.5), and the integration of DLC with various decoding methods (Appendix A.6).
+We primarily validate the effectiveness of the adaptive factor $\lambda_{t}$.
+The Appendix provides a large number of additional experiments, including comprehensive ablation studies on key components and hyperparameters.
+These experiments also demonstrate DLC’s robustness across different CLIP models, its broader applicability to VQA tasks, and its scalability to larger 13B architectures.
 Effectiveness of the adaptive factor in $\lambda_{t}$.
 To demonstrate the effectiveness of the adaptive factor $(1-\bar{B}_{t})^{2}$ in dynamic guidance strength $\lambda_{t}$, as formulated in Eq.
-8, we configure models with and without this factor to generate responses on the COCO [28] dataset.
+8, we configure models with and without this factor to generate responses on the COCO (Lin et al., 2014) dataset.
 The correctness ($C$) and detailedness ($D$) of these responses are evaluated using GPT-4o.
@@
 Section A.5 analyzes the impact of two key hyperparameters on performance: the modulation strength $\alpha$ and the historical context window size $N$.
+Section A.6 analyzes the robustness of DLC across different CLIP model variants.
+Section A.7 presents the extension of DLC to Visual Question Answering tasks, demonstrating its applicability beyond open-ended generation.
+Section A.8 provides comprehensive ablation studies validating the importance of each component in our framework.
+Section A.9 demonstrates DLC’s effectiveness on larger 13B-parameter models, confirming its scalability beyond 7B architectures.
 Section A.10 demonstrates the integration of DLC with various decoding strategies, including beam search, greedy search, top-k sampling, and temperature scaling.
@@
 A.1 More backgrounds about LVLMs
-The development of LVLMs has evolved from BERT-based [14, 33, 38] cross-modal alignment to LLM-powered [1, 7, 12, 26, 30, 56] architectures.
-Building upon foundational text-image alignment techniques like CLIP [35] and BLIP [25], contemporary LVLMs integrate visual encoders with LLM decoders through distinct optimization strategies.
-For instance, LLaVA [30] employs lightweight adapters combined with instruction tuning, while InstructBLIP [12] and MiniGPT-4 [56] utilize Q-Former modules for visual token compression.
+The development of LVLMs has evolved from BERT-based [Devlin et al., 2019, Lu et al., 2019, Tan and Bansal, 2019] cross-modal alignment to LLM-powered [Alayrac et al., 2022, Chen et al., 2023, Dai et al., 2023, Li et al., 2023b, Liu et al., 2023, Zhu et al., 2023] architectures.
+Building upon foundational text-image alignment techniques like CLIP [Radford et al., 2021] and BLIP [Li et al., 2022], contemporary LVLMs integrate visual encoders with LLM decoders through distinct optimization strategies.
+For instance, LLaVA [Liu et al., 2023] employs lightweight adapters combined with instruction tuning, while InstructBLIP [Dai et al., 2023] and MiniGPT-4 [Zhu et al., 2023] utilize Q-Former modules for visual token compression.
 Despite these innovations, significant challenges persist, particularly severe hallucination issues arising from misaligned multimodal representations.
@@
 2:Generated text sequence $y$
-3:$y\leftarrow x$triangleright Initialize output sequence with input prompt
-4:$\bar{B}\leftarrow 0$triangleright Initialize historical baseline
+[…]
+[…]
 […]
@@
 These details are essential for reproducing our results and understanding the practical aspects of DLC.
-We compute all visual alignment scores using the SigLIP model [51].
-The historical baseline $\bar{B}_{t}$ is calculated over a sliding window of $N=8$ past context scores, following a 3-step warm-up period.
-For the top-k candidates $(k=50)$ at each step, the combined score $S_{t,i}^{\text{COMB}}$ is computed.
-The adaptive modulation employs a maximum intervention strength $\alpha=3$, and the standard sigmoid function $\sigma(\cdot)$ for logit adjustment.
-Our experiments were run on a single A100 GPU.
+[…]
 A.4 Further Details on Benchmarks
@@
 This section provides additional details about these benchmarks, including dataset characteristics, evaluation protocols, and metrics interpretation.
-(1) CHAIR [36] quantifies object hallucination in image captions by measuring the ratio of objects mentioned in captions but absent from ground-truth labels.
+(1) CHAIR Rohrbach et al.
+[2018] quantifies object hallucination in image captions by measuring the ratio of objects mentioned in captions but absent from ground-truth labels.
 It includes two variants: CHAIRI ($C_{I}$), calculating hallucinated objects versus all generated objects, and CHAIRS ($C_{S}$), assessing sentences with hallucinations versus total sentences.
@@
 Lower scores indicate better hallucination suppression.
-(2) GPT-4 Assisted Evaluations [54] function as a crucial complement to conventional assessment methodologies.
+(2) GPT-4 Assisted Evaluations Zhao et al.
+[2023] function as a crucial complement to conventional assessment methodologies.
 While traditional metrics such as CHAIR effectively quantify hallucinations at the object-existence level, they demonstrate inadequacy in detecting more fine-grained hallucination types, including positional, relational, and attribute-based inaccuracies.
-To address this limitation, GPT-4 assisted evaluation employs detailed object-level descriptions from the Visual Genome (VG) [20] dataset as ground-truth reference, leveraging GPT-4’s sophisticated capabilities for precise hallucination detection and classification.
+To address this limitation, GPT-4 assisted evaluation employs detailed object-level descriptions from the Visual Genome (VG) Krishna et al.
+[2017] dataset as ground-truth reference, leveraging GPT-4’s sophisticated capabilities for precise hallucination detection and classification.
 The complete prompt template is provided in the Appendix A.12.
 (3) GPT-4o assisted evaluations [^1] We utilize GPT-4o for assisted evaluation, as it offers enhanced capabilities over previous vision-capable models like GPT-4V, which is also being phased out by OpenAI.
+Yin et al.
 [2024] further conduct a comprehensive analysis of hallucinations and text quality in open-ended generation tasks.
-Harnessing GPT-4o’s sophisticated capabilities, [47] established a numerical assessment protocol utilizing a 0-10 scale, which evaluates two essential dimensions: Accuracy-the extent to which responses align with image content without fabrication, and Detailedness-the degree of informational richness present in the responses.
-(4) LLaVA-Bench-in-the-Wild [30] is designed to evaluate LVLMs’ proficiency in addressing complex tasks and their adaptability to novel domains.
+Harnessing GPT-4o’s sophisticated capabilities, [Yin et al., 2024] established a numerical assessment protocol utilizing a 0-10 scale, which evaluates two essential dimensions: Accuracy-the extent to which responses align with image content without fabrication, and Detailedness-the degree of informational richness present in the responses.
+(4) LLaVA-Bench-in-the-Wild Liu et al.
+[2023] is designed to evaluate LVLMs’ proficiency in addressing complex tasks and their adaptability to novel domains.
 We perform targeted case studies on this dataset to demonstrate the efficacy of our proposed DLC.
@@
 In this section, we conduct a comprehensive sensitivity analysis to understand how these parameters affect the model’s hallucination mitigation capability and generation quality.
-Experiments were conducted using a range of values for these parameters across three distinct LVLMs, with performance assessed by the CHAIR benchmark [36] (max new tokens: 512 for all ablation studies).
+Experiments were conducted using a range of values for these parameters across three distinct LVLMs, with performance assessed by the CHAIR benchmark [Rohrbach et al., 2018] (max new tokens: 512 for all ablation studies).
 Table 4: Impact of Modulation Strengths $\alpha$ on CHAIR Metrics for Different LVLMs and Decoding Strategies
@@
 While default values are used for general demonstration in the main body of this paper, the ablation studies presented here clearly indicate that optimal performance is achieved when these parameters are carefully tuned according to the specific LVLM and the selected decoding strategy.
-A.6 DLC Performance with Various Decoding Strategies
+A.6 Robustness Analysis Across Different CLIP Models
+To analyze the robustness of our DLC method and its dependence on the underlying CLIP model, we conducted comprehensive experiments using different CLIP backbone architectures.
+This analysis addresses concerns about the method’s reliance on a specific vision-language alignment model.
+Table 6 presents the performance of DLC when integrated with various CLIP variants, including the original CLIP models, SigLIP, and the recent FG-CLIP model.
+Table 6: Robustness Analysis Across Different CLIP Models. CHAIRS and CHAIRI scores (lower is better) for DLC when using different CLIP backbones. Best results are highlighted in bold.
+CLIP Backbone LLaVA-1.5 InstructBLIP
+[…]
+CLIP-ViT-Large-Patch14 40.4 11.4 51.8 14.7
+CLIP-ViT-Large-Patch14-336 39.7 10.6 51.2 14.4
+SigLIP-SO400M-Patch14-384 39.4 10.9 51.0 14.3
+FG-CLIP-Large [Xie et al., 2025] 37.8 10.4 49.8 13.9
+The results demonstrate that while stronger CLIP models generally lead to better performance, DLC maintains consistent improvements across all tested architectures.
+This robustness indicates that our method is not brittlely tied to a specific CLIP variant, but rather benefits from the general principle of visual-textual alignment assessment.
+The superior performance with FG-CLIP, which specializes in fine-grained visual-textual alignment, particularly validates our approach’s potential for further enhancement as CLIP models continue to evolve.
+A.7 Extension to Visual Question Answering Tasks
+While DLC was initially designed for open-ended descriptive generation tasks, we have successfully extended it to Visual Question Answering (VQA) domains through simple modifications.
+By treating the user’s query content as already generated context, DLC demonstrates excellent performance on VQA tasks that require semantic-level reasoning.
+We validated this extension on the popular POPE benchmark [Li et al., 2023c], which evaluates object existence in images through yes/no questions across three different setups: random, popular, and adversarial.
+The results are presented in Table 7.
+The results demonstrate the effectiveness of DLC, validated by its successful extension from open-ended generation to structured VQA scenarios.
+Table 7: Performance on POPE Benchmark.
+Results show DLC’s effectiveness when extended to VQA tasks.
+Accuracy (Acc) and F1 scores are reported for different setups.
+Best results are highlighted in bold.
+Method Setup LLaVA-1.5 InstructBLIP
+[…]
+Sample Random 83.90 84.23 83.13 83.43
+ICD 85.33 85.48 82.46 83.68
+VCD 82.50 83.20 83.60 83.69
+SID 87.56 87.10 84.36 84.43
+DLC+SigLIP (Ours) 87.83 88.14 88.93 88.81
+DLC+FG-CLIP (Ours) 88.87 88.88 89.53 89.32
+Sample Popular 80.93 81.85 77.20 78.78
+ICD 82.16 82.80 74.20 77.66
+VCD 79.86 81.13 77.13 78.68
+SID 84.16 84.11 77.10 78.98
+DLC+SigLIP (Ours) 84.37 85.23 82.73 83.69
+DLC+FG-CLIP (Ours) 85.33 85.84 82.53 83.42
+Sample Adversarial 75.57 77.83 74.80 77.02
+ICD 76.33 78.39 72.03 76.15
+VCD 75.10 77.66 75.06 76.99
+SID 79.60 80.43 76.03 78.07
+DLC+SigLIP (Ours) 77.00 79.68 79.03 80.82
+DLC+FG-CLIP (Ours) 78.70 80.68 79.57 81.10
+A.8 Comprehensive Component Ablation Studies
+To thoroughly validate the importance of each component in our proposed DLC framework, we conducted comprehensive ablation studies.
+We systematically evaluated the impact of removing individual components: Current Contextual Token Alignment (CCTA), Isolated Token Alignment (ITA), and Relative Visual Advantage (RVA).
+The ablation experiments were conducted using GPT-4o assisted evaluation on the COCO dataset, with models generating detailed descriptions for 500 randomly selected images.
+Performance is assessed on two key dimensions: Correctness (accuracy of visual descriptions) and Detailedness (richness of information).
+Table 8: Comprehensive Ablation Study Results. GPT-4o evaluation scores (0-10 scale, higher is better) demonstrating the importance of each DLC component. Best results are highlighted in bold.
+Settings LLaVA-1.5 InstructBLIP
+[…]
+w/o CCTA 6.22 6.65 5.79 6.20
+w/ CCTA 7.00 6.72 6.60 6.46
+w/o ITA 6.22 6.61 5.84 6.30
+w/ ITA 6.97 6.72 6.63 6.52
+w/o RVA 6.35 6.51 5.83 6.25
+w/ RVA 7.00 6.69 6.71 6.60
+The results clearly demonstrate that each component contributes significantly to the overall performance of DLC.
+A.9 Scalability to Larger Model Architectures
+To validate the effectiveness of our method on larger-scale models, we conducted comprehensive experiments on 13B-parameter architectures, specifically evaluating DLC on LLaVA-1.5-13B and InstructBLIP-13B using both the CHAIR benchmark and GPT-4o assisted evaluation.
+A.9.1 CHAIR Benchmark Results on 13B Models
+Table 9 presents the CHAIR evaluation results for 13B models across different maximum token settings.
+The results consistently demonstrate that DLC provides substantial improvements over nucleus sampling baselines on both model architectures.
+The results show remarkable improvements across both token length settings
+Table 9: CHAIR Evaluation Results on 13B Models.
+Comparison of nucleus sampling baseline and DLC on larger model architectures.
+Lower CHAIR scores indicate better performance (fewer hallucinations).
+Best results are highlighted in bold.
+Methods Max Token LLaVA-1.5-13B InstructBLIP-13B
+[…]
+Nucleus Sampling 64 23.20 8.16 27.80 10.22
+DLC (Ours) 64 18.20 5.53 24.20 7.75
+Nucleus Sampling 512 53.20 15.07 64.00 19.57
+DLC (Ours) 512 33.00 8.80 55.60 15.37
+A.9.2 GPT-4o Assisted Evaluation on 13B Models
+To complement the CHAIR evaluation, we conducted GPT-4o assisted evaluation focusing on correctness and detailedness of generated descriptions. The results are presented in Table 10.
+The GPT-4o evaluation demonstrates substantial improvements in both correctness and detailedness.
+Table 10: GPT-4o Evaluation Results on 13B Models. Performance comparison on correctness and detailedness metrics. Best results are highlighted in bold.
+Settings LLaVA-1.5-13B InstructBLIP-13B
+[…]
+Nucleus Sampling 5.25 6.05 4.26 5.06
+DLC (Ours) 7.75 6.77 7.00 6.36
+A.10 DLC Performance with Various Decoding Strategies
 Our proposed DLC demonstrates compatibility with various mainstream decoding strategies.
-To comprehensively evaluate its effectiveness, in addition to its performance with common strategies such as nucleus sampling and greedy sampling, we conducted further experiments on the LLaVA-1.5 [30] model using the MSCOCO dataset [28] and the CHAIR benchmark [36].
+To comprehensively evaluate its effectiveness, in addition to its performance with common strategies such as nucleus sampling and greedy sampling, we conducted further experiments on the LLaVA-1.5 [Liu et al., 2023] model using the MSCOCO dataset [Lin et al., 2014] and the CHAIR benchmark [Rohrbach et al., 2018].
 These experiments explored other decoding methods and parameters, specifically: Top-p sampling (p=0.9), Top-k sampling (k=50), and Top-k sampling (k=50) with varying temperatures (t=1.5 and t=0.8).
@@
 Figure 8: Results of different decoding strategies.
-A.7 Discussion of Methodological Limitations and Social impacts.
+A.11 Discussion of Methodological Limitations and Social impacts.
 In this section, we delineate two limitations inherent to the proposed DLC method:
-Scope of Applicability: The DLC method is primarily optimized for open-ended, descriptive text generation scenarios, aiming to mitigate semantic drift that may occur during the generation of lengthy texts by dynamically calibrating the probabilities of candidate tokens.
-Consequently, its advantages may not be fully realized in contexts requiring minimalistic answers, such as "yes/no" judgments (e.g., as encountered in tasks like the POPE benchmark [27]).
-The core mechanism of DLC relies on the generated textual context for dynamic adjustments; for extremely short answers, such as single "yes" or "no" responses, the lack of sufficient contextual information for effective logits calibration might constrain its performance.
-To address this limitation, future work could explore transforming such concise answers (e.g., "yes" or "no") into complete declarative sentences.
-For instance, for object presence verification, sentences like "There is an <object> in the image" versus "There is not an <object> in the image" could be constructed.
-Subsequently, CLIP could be employed to compute the semantic similarity between these statements and the input image, thereby assisting the model in selecting a more accurate response.
-This direction warrants further exploration and validation.
-Dependency on CLIP Model Performance: The efficacy of DLC is intrinsically linked to the performance of the employed CLIP model.
-The CLIP model is tasked with assessing the semantic similarity between images and text, which serves as the foundational input for DLC’s logits calibration process.
-Therefore, if the CLIP model itself exhibits deficiencies or biases in accurately capturing the semantic associations for specific image-text pairs-stemming, for example, from biases in its training data or an insufficient understanding of certain complex, fine-grained visual concepts-this can directly impact the precision and overall effectiveness of the DLC calibration.
-Theoretically, leveraging more advanced models with superior vision-text alignment capabilities is anticipated to further enhance the performance ceiling of DLC.
+Limited Applicability to Spatial and Coordinate-Level Tasks: While our extension to VQA tasks shows promising results on object existence verification (POPE benchmark [Li et al., 2023c]), DLC faces inherent limitations with tasks requiring precise spatial reasoning or coordinate-level grounding.
+Our method relies on CLIP’s global semantic assessment, which lacks sensitivity to fine-grained spatial details or numerical coordinates.
+This limitation affects tasks such as: (1) object counting and spatial relationship reasoning (e.g., "How many red cars are to the left of the blue building?"), (2) precise coordinate grounding tasks like object detection with bounding box generation, and (3) referring expression comprehension requiring exact localization.
+Dependence on Visual Content Quality: DLC’s effectiveness is inherently tied to the quality and clarity of the input visual content.
+For images with poor quality, ambiguous visual information, the method’s ability to provide reliable calibration signals may be compromised.
+In such cases, the visual alignment scores from CLIP may not provide sufficiently discriminative information to guide token selection effectively, potentially limiting the method’s benefits.
 Social impacts.
@@
 DLC not only provides an effective and efficient solution to a key challenge but also serves as an inspiration for the research community to explore new pathways for improving visual grounding and tackling dynamic vision-language alignment, fostering the development of more robust and trustworthy LVLMs.
-A.8 Extended Comparative Results on LLaVA-Bench
-In this section, we present additional comparative results of our DLC approach against other methods on the LLaVA-Bench [30].
+A.12 Prompts for GPT-4 and GPT-4o Evaluation
+Figure 9: Prompts of GPT-4 for evaluations.
+Figure 10: Prompts of GPT-4o for evaluations.
+A.13 Extended Comparative Results on LLaVA-Bench
+In this section, we present additional comparative results of our DLC approach against other methods on the LLaVA-Bench [Liu et al., 2023].
 These extended evaluations provide a more comprehensive understanding of our method’s performance across various question types and scenarios.
 Figure 11: Case Study: LLaVA-1.5 on LLaVA-Bench-in-the-Wild (Hallucinations Marked Red).
-A.9 Prompts for GPT-4 and GPT-4o Evaluation
-Figure 10: Prompts of GPT-4 for evaluations.
-Figure 11: Prompts of GPT-4o for evaluations.